In a P-Type semiconductor, the Fermi level in 0.3 eV above the valance band at a room temperature of 300 oK. Determine the new position of the Fermi level for temperature of a (350 oK) and (b) 400 oK.


1
Expert's answer
2021-07-04T19:54:02-0400

"E_F=E_V+KT\ln(\frac{N_v}{N_A})\\
E_F-E_V=0.3eV\\
T=300\degree{K}\\
E_F-E_V=KT\ln(\frac{N_v}{N_A})\\
0.3=300K\ln(\frac{N_v}{N_A})\\
K\ln(\frac{N_c}{N_A})=\frac{0.3}{300}\\
K\ln(\frac{N_c}{N_A})=0.001\\
(a)\\
T=350\degree{K}\\
E_F-E_V=KT\ln(\frac{N_v}{N_A})\\
E_F-E_V=350K\ln(\frac{N_v}{N_A})\\
E_F-E_V=350\times 0.001\\
E_F-E_V=0.35eV\\
(b)\\
T=400\degree{K}\\
E_F-E_V=KT\ln(\frac{N_v}{N_A})\\
E_F-E_V=400K\ln(\frac{N_v}{N_A})\\
E_F-E_V=0.40eV"